# Towards universal property prediction in Cartesian space: TACE is all you need

### v1

@@ 2.12 Out-of-Domain — 3BPA
The root-mean-squared errors (RMSE) on both energy and forces for several benchmark models are reported in Table 8. We find that TACE exhibits strong extrapolation capabilities, especially in force prediction, where it reaches exceptionally high accuracy. Moreover, the 3BPA dataset is employed in our study to investigate the differences between coupled and uncoupled feature representations in Eqn. 12. We observe that coupled features consistently achieve higher accuracy than the uncoupled features on force predictions across all test sets with different channels. TACE already reaches SOAT at 64 channels, whereas MACE was trained with 256 channels. This difference may be attributed to MACE’s use of uncoupled features. Nonetheless, this result indicates that, compared to spherical harmonic-based models, Cartesian models may achieve higher accuracy with fewer channels when using coupled channels and if lower-weights can be effectively utilized, the channel can be further reduced.
For universal machine learning interatomic potentials (UMLIPs), which typically involve multiple chemical elements, many current models include as many as 89 elements. To achieve nearly highest accuracy, for example, the coupled model configured with $c=64$ and $\mathrm{L_{max}}=0,1,2$ comprises approximately 21M, 37M, and 55M parameters, respectively, whereas the uncoupled model has only about 0.8M, 1.1M, and 1.4M parameters. Although the coupled model contains significantly more parameters as the number of elements increases, the computational cost is primarily determined by $\mathrm{L_{max}}$ and these element dependent weights are just use in single MLP. Therefore, the use of coupled features has a negligible impact on computational efficiency. If the number of parameters is considered excessive, it can be effectively reduced by setting the hidden channel to 32, which decreases the total parameter count from 55M to 14M.
Table 8: RMSE for energy (E, meV) and forces (F, meV/Å) for the 3BPA dataset. The best-performing results are both bolded and underlined.22footnotetext: All models were trained at $T=300\,\mathrm{K}$. Results for all models except TACE are taken from Ref. [14]. Here, both TACE models achieve excellent accuracy. However, note that the use of “all” is merely intended to demonstrate its potential applicability, and does not represent the optimal way of employing it. We only show the performance of “all” here, while all other benchmarks are based on “max”.
GAP[55] ACE[24] CACE[14] Allegro[8] NequIP[7] MACE[9] $\mathbf{TMP_{all}}$ $\mathbf{TACE_{max}}$ $\mathbf{TACE_{all}}$
300K E 22.8 7.1 6.3 3.8 3.3 3.0 3.16 2.69 2.87
F 87.3 27.1 21.4 13.0 10.8 8.8 10.62 8.68 9.84
600K E 61.4 24.0 18.0 12.1 11.2 9.7 10.99 9.09 9.20
F 151.9 64.3 45.2 29.2 26.4 21.8 26.20 19.88 22.59
1200K E 166.8 85.3 58.0 42.6 38.5 29.8 42.38 23.42 31.90
F 305.5 187.0 139.6 83.0 76.2 62.0 84.71 54.26 63.10
dihedral E - - - - 23.2 7.8 13.82 11.00 13.82
F - - - - 23.1 16.5 18.37 16.45 18.99
2.13 Advantages and limitations of the ACE framework
We note that NequIP [7] was the earliest high-accuracy equivariant neural networks. Since then, with the rapid development of equivariant architectures, numerous advanced frameworks have emerged. Among these, ACE [24] and its variants—namely MACE [9], GRACE [26], CACE [14], and the TACE model introduced in this work are particularly representative. Typically, these ACE-based models employ only two layers, that is, a single message-passing step, which is already sufficient to reach the accuracy of non-ACE models with 3 to 5 layers. Adding further layers generally provides limited benefit within the ACE framework. This characteristic brings both advantages and disadvantages for practical applications.
For instance, compared to NequIP, MACE removes gating operations and introduces self-product terms, making it approximately 3 to 4 times faster [9]. Owing to the strong inductive bias of ACE, MACE can achieve excellent accuracy for most systems. However, the reduced number of layers limits the effective receptive field. For example, in the Al-doped Au2 MgO(001) system [19], MACE exhibits poorer performance at T=2 without long-range corrections. In principle, increasing the number of layers could enlarge the receptive field, but this would inevitably compromise efficiency. A similar trade-off also applies to TACE.
With the rapid development of datasets, more and more researchers are realizing the importance of non-equilibrium structures for model training. A representative dataset is OMat24 [66], on which numerous model architectures have been trained, demonstrating both good stability and generalization performance. The RECIO [67] sampling method is a system-independent, nonequilibrium-structure-dominated selection strategy. In its original publication, RECIO combined with NequIP was used to construct a five-element potential for PdAgCHO, achieving wide applicability in various catalytic reactions without system-specific sampling.
However, when we attempted to train TACE and MACE using the dataset provided in that work, we observed a consistent issue. During training, the MAE decreased as expected, but after the validation set error reached a certain point, the RMSE suddenly began to increase steadily, even while the MAE continued to decrease. Eventually, both errors started to rise, preventing effective training. We experimented with training strategies that avoided reliance on validation-based learning rate schedulers, as well as with different loss functions and a series of other techniques. We found that stable training was only possible under very specific hyperparameter settings, which were extremely restrictive. Further investigation showed that fine-tuning MACE-MPA and MACE-OMAT [68] on this dataset led to different behaviors: MACE-OMAT could be trained robustly under all conditions, while MACE-MPA was more sensitive and tended to diverge when fine-tuned with a small number of data points.
Since this instability was observed in both TACE and MACE, we attribute the problem to an inherent conflict with ACE and RECIO. Although RECIO is sufficiently general to be applied across diverse systems, the generated structures now lack enough physical constraints, which leads to conflicts with the strong inductive biases of ACE-based frameworks. As a solution, we propose incorporating LES as a plugin. LES serves as a long-range correction term, which is particularly well suited for ACE models with relatively small receptive fields. Moreover, its flexible choice of Coulombic, London dispersion or other forms introduces appropriate physical constraints on energy and forces, thereby achieving excellent compatibility with TACE and related models. Next, we will use the PdAgCHO dataset generated by the RECIO method to test various reactions. The dataset contains a total of 116,516 structures, and when the cutoff is set to 5 Å, the average number of neighboring atoms is 23.8. The following examples successfully demonstrates that TACE can effectively learn from the complex datasets generated by the RECIO method [67]. Although this dataset only contains bulk structures, the trained MLIP can still accurately and efficiently simulate various slabs. We employed the Atomic Simulation Environment package [69] to perform the following structure optimizations and transition state (TS) searches.
2.13.1 Adsorption of PTCDA on Ag(111)
We also quantitatively analyzed the interatomic distances within the adsorption complex like ref [67]. The average distance comparison with experiment [70], DFT [71], NequIP [7, 67], UMLIPs (MACE-OMAT [68], OrbV3-onservative-inf-omat [72], 7Net-omat [73]) and fine-tuned MACE-OMAT are shown in Table 9. It should be noted that only the data among UMLIP models are comparable, whereas the TACE results should be compared with DFT and NequIP. Our original intention was to compare the UMLIP of the two SOATs. However, in the final selection process, we only fine-tuned MACE-OMAT because the conservative version of OrbV3 had some issues during our fine-tuning, and the direct version performed poorly. Even after fine-tuning MACE-OMAT, the results were still not satisfactory. Our preliminary conclusion is that the conflict between ACE and RECIO remains, so although there are no anomalies in the error table, the effect of fine-tuning was indeed not good.
Table 9: Average distances of main chain carbon, carboxylic oxygen, and anhydride oxygen atoms from the Ag(111) surface (in Å).22footnotetext: The benchmark result of the TACE model was obtained with a batch size of 16 at the 216th epoch and MAE are compared against DFT.
EXP [70] DFT [71] OrbV3 7Net MACE-OMAT MACE-OMAT-fine-tuned NequIP [67] TACE-LES
main chain carbon 2.86 2.97 3.38 3.22 3.21 3.00 2.98 2.97
carboxylic oxygen 2.66 2.67 3.07 3.03 3.40 2.72 2.68 2.69
anhydride oxygen 2.98 2.85 3.18 3.15 3.51 2.99 2.92 2.86
MAE - - 0 .38 0.30 0.54 0.07 0.03 0.01
2.13.2 CO oxidation on various PdAg surfaces
CO oxidation, a widely employed model reaction in heterogeneous catalysis, has played a pivotal role in advancing our understanding of fundamental catalytic mechanisms. To evaluate the ability of TACE to distinguish and accurately capture different surface facets, we investigated the reactivity across a range of PdAg surfaces, including low-index surfaces (Pd(100), Pd(110), Pd(111)), specialized surfaces (twin boundary, Pd-TB; nanorod structure, Pd-nanorod), as well as various Pd configurations on Ag(111) (single-atom Pd/Ag(111) SAC, Pd adatom on Ag(111), Pd clusters, and Pd lines on Ag(111)). We verified the reliability of the predicted TS geometries by confirming their structural correctness and ensuring that each valid TS exhibited exactly one imaginary frequency (as the Hessians were available, frequencies could be obtained either by finite-difference calculations or by diagonalization after coordinate transformation, with the latter introducing less numerical noise). The corresponding energy barriers are summarized in Table 10.
Table 10: Energy barriers (eV) for CO oxidation on various PdAg surfaces. Failures in searching TS are indicated by ×.22footnotetext: The benchmark result of the TACE model was obtained with a batch size of 16 at the 216th epoch and MAE are compared against DFT.
System DFT [67] OrbV3-con-inf MACE-OMAT MACE-OMAT-fine-tuned NequIP [67] TACE-LES
Pd100 0.63 0.60 0.66 0.69 0.67 0.65
Pd110 0.69 0.64 0.69 0.70 0.79 0.67
Pd111 1.39 × × × 1.35 1.29
Pd1/Ag(111) 0.28 0.27 × × 0.38 0.29
Pd4/Ag(111) 0.23 × 0.28 × 0.21 0.20
Pd9/Ag(111) 0.87 × × × 0.73 0.87
Pd13/Ag(111) 0.74 × 0.36 0.45 0.60 0.54
Pd20/Ag(111) 0.90 0.69 0.89 0.92 0.99 0.91
Pd-TB 0.86 0.71 1.04 1.17 0.91 1.12
Pd-cube 1.28 1.01 1.17 1.23 1.18 1.23
Pd-S/Ag(111) 0.42 0.46 × × 0.42 0.38
Pd-line/Ag(111) 0.55 × × 0.51 0.58 0.51
MAE - 0.11 0.11 0.11 0.07 0.07
2.13.3 Acetylene hydrogenation on Pd(111), H covered Pd(111) and AgPd alloys
To validate the reliability of TACE for surface coverage eaction pathways, we calculated for acetylene hydrogenation on both clean Pd(111) and hydrogen-covered Pd(111) surfaces. The results are shown in Table 11.
Table 11: Energy barriers (eV) for various Pd(111) and PdAg surfaces with different adsorbates. Failures in searching TS are indicated by ×.22footnotetext: The benchmark result of the TACE model was obtained with a batch size of 16 at the 216th epoch and MAE are compared against DFT.
System DFT [67] OrbV3-con-inf MACE-OMAT MACE-OMAT-finetuned NequIP [67] TACE-LES
Pd(111)-C2H3 0.87 0.72 0.74 0.78 0.73 0.79
Pd(111)-C2H4 0.79 0.70 0.76 0.72 0.74 0.71
Pd(111)-C2H5 0.79 0.67 0.61 0.74 0.73 0.85
Pd(111)-C2H6 0.71 × 0.50 0.53 0.76 0.70
1 ML-Pd(111)-C2H3 0.67 × 0.72 × 0.74 0.36
1 ML-Pd(111)-C2H4 0.69 0.65 0.65 0.64 0.62 0.57
1 ML-Pd(111)-C2H5 0.90 × 0.60 0.72 0.64 0.75
1 ML-Pd(111)-C2H6 0.45 0.45 0.39 0.45 0.42 0.48
0.25 ML-PdAg3(111)-C2H3 0.58 × 0.61 0.60 0.51 0.63
0.25 ML-PdAg3(111)-C2H4 0.37 0.59 0.52 0.47 0.44 0.43
0.25 ML-PdAg3(111)-C2H5 0.42 0.67 0.61 0.68 0.71 0.62
0.25 ML-PdAg3(111)-C2H6 0.17 0.33 × 0.25 0.35 0.32
1 ML-PdAg(111)-C2H3 0.58 0.52 0.52 0.52 0.48 0.49
1 ML-PdAg(111)-C2H4 0.36 × 0.67 0.45 0.43 0.35
1 ML-PdAg(111)-C2H5 0.79 0.78 0.67 0.72 0.46 0.70
1 ML-PdAg(111)-C2H6 0.36 0.44 0.41 0.39 0.46 0.45
MAE - 0.11 0.13 0.09 0.12 0.10
2.13.4 Baker reaction system
We selected eight Baker reaction systems [74], representing a variety of reaction types, as benchmark examples for small-molecule transition state searches. The results are summarized in Table 12.
Table 12: Baker reaction system energies (eV). Failures are indicated by ×.22footnotetext: The benchmark result of the TACE model was obtained with a batch size of 16 at the 216th epoch and MAE are compared against DFT.
System DFT [67] OrbV3-con-inf MACE-OMAT MACE-OMAT-fine-tuned NequIP [67] TACE-LES
HCCH $\leftrightarrow$ CCH2 1.98 × × × 2.09 2.06
H2CO $\leftrightarrow$ H2 + CO 2.83 2.83 × 3.11 3.05 3.00
bicyclo[1,1,0] $\leftrightarrow$ butane TS1 2.66 × 2.33 3.03 2.75 2.87
butadiene + ethylene $\leftrightarrow$ cyclohexene 0.46 0.58 × 0.84 0.67 0.65
trans-butadiene $\leftrightarrow$ cis-butadiene 0.32 0.28 0.61 0.28 0.37 0.24
vinyl alcohol $\leftrightarrow$ acetaldehyde 2.25 2.37 2.11 2.37 2.02 2.23
claisen rearrangement 1.65 1.88 2.02 2.211 1.67 1.90
rotational TS in acrolein 0.38 0.39 0.53 0.40 0.38 0.34
MAE - 0.09 0.26 0.25 0.12 0.13
@@ 3 Discussion
At present, Cartesian models based on multi-layer architectures (T >> 2) have already surpassed spherical-harmonics models in computational speed at low ranks, without any loss in accuracy. However, in terms of accuracy, we believe that the ACE framework has greater advantages. For Cartesian models based on ACE (typically with T = 2), such as the one presented in this work, their computational speed has not yet surpassed that of spherical-harmonics models, mainly due to the relatively inefficient implementation of Cartesian self-contraction in the current framework. Our work has already demonstrated accuracy that is comparable to or surpasses that of SOAT models, and its performance is expected to be further enhanced through a thorough investigation of low-weight transformations, potentially enabling Cartesian models to fully surpass spherical-harmonics models at low ranks ($\leq 4$).
@@ 4 Conclusion
In this work, we present TACE and TMP, both of which are E(3)-equivariant message-passing frameworks built upon higher-order Cartesian tensors and operating directly in Cartesian space. TACE accommodates tensors of arbitrary rank and both reducible or irreducible Cartesian tensors. The framework can consistently extract the corresponding irreducible components of different weight, allowing us to freely extract and manipulate them. Furthermore, the model is capable of directly predicting tensorial properties, and through the use of a universal embedding, we can flexibly incorporate and control various factors influencing the predicted physical quantities. We believe that this represents the future trend and methodology in the development of MLIPs, and with universal embedding both invariants and equivariants can be handled in a unified manner. Although we only demonstrate external fields, magnetic moments, charges, and basis sets here, the framework can be readily extended. In future work, we will focus on pursuing more efficient Cartesian tensor contractions and backend optimizations, while also providing interfaces to various software packages, thereby enabling more practical and direct applications of the approaches presented here.
@@ 5 Methods
All experiments were run with the TACE software available at https://github.com/xvzemin/tace. In addition, the Pytorch version was 2.6.0+cu124, PyTorch Lightning under version 2.5.2, and Python under version 3.12.11. For other models, we use orb-models 0.5.2, mace-torch 0.3.12, sevenn 0.11.1.dev2 and allegro-pol 0.0.3.
Datasets
rMD17: For rMD17 dataset [64], we employ a cutoff radius of 5 Å. The rMD17 dataset comprises ten distinct molecules, each containing 100,000 structures, except for azobenzene which includes 99,988 structures. Depending on the experimental setup, the training set contains either 50 or 950 structures, the validation set is consistently fixed at 50 structures, and all remaining structures constitute the test set.
@@
Water: For the liquid water dataset [37], a cutoff of 6 Å is applied. Following the protocol used in CAMP [15], the dataset is split into 1,432 configurations for training, 79 for validation, and 81 for testing. The dipole moment ($\mu$) and polarizability ($\alpha$) for various water systems were taken from Ref. [42]. Values of $\mu$ and $\alpha$ for molecules such as $\mathrm{H_{2}O}$, $(\mathrm{H_{2}O})_{2}$, and $\mathrm{H_{5}O_{2}}^{+}$ were calculated using the coupled cluster singles and doubles (CCSD) method with the d-aug-cc-pVTZ basis set. For liquid water, the dataset was generated using the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional together with ultra-soft pseudopotentials (USPPs). A cutoff of 6 Å was employed for training the dipole and polarizability models, except for $(\mathrm{H_{2}O})_{2}$, for which a cutoff of 10 Å was used. The training and validation splits were either 500:500 or 700:300. The dataset employed for IR and Raman simulations comprises 1,888 structures calculated using the strongly constrained and appropriately normed (SCAN) functional. Our model attains a RMSE of 2.43meV/atom for energy, 30.97meV/Å for forces, and 1.00 meV/atom for virials. The train-test split follows the TNEP scheme [50], with 1,510 structures used for training and 378 for testing, and an additional 10% of the training set reserved for validation.
@@
Charges: The datasets for different charge states and charge-transfer systems [35] were generated using the PBE functional with a light computational setting. Spin-polarized calculations were performed for the Au2-MgO, NaCl, and Ag3 systems. The Ag, NaCl, carbon-chain, and Au2-MgO datasets contain 11,013, 5,000, 10,019, and 5,000 structures, respectively. For each dataset, 10% of the configurations were used as the validation set, and all reported errors are evaluated on the validation set. For the charge embedding benchmarks of the Ag, NaCl, carbon-chain, and Au2-MgO systems, the cutoff radii are 5.29, 5.29, 4.23, and 5.0 Å, respectively. When using charges together with energy and forces as training labels, the corresponding cutoff values are 6.0, 6.0, 4.23, and 6.0 Å, respectively.
@@
PdAgCHO: PdAgCHO dataset was generated using the RECIO method [67] and contains a total of 116,516 structures. For training, 100,000 structures were randomly selected, while the remaining configurations were used for validation. All structures were computed at the PBE level. The training of UMLIPs, fine-tuning, and TACE models was carried out by us with a uniform cutoff of 6 Å. It should be noted that the NequIP [7] model was not trained by us, and the dataset partitioning scheme used in NequIP differs from ours.
@@
IR and Raman spectra
In this work, the IR and Raman spectra were obtained by applying a Fourier transform to the corresponding autocorrelation functions (ACF), thereby converting from the time domain to the frequency domain, following the formulation outlined in refs. [50, 43]. The IR intensity, as well as the isotropic and anisotropic Raman intensities, are expressed as:
$I(\omega)\propto\omega^{2}\int_{-\infty}^{+\infty}e^{-i\omega t}\Big\langle\boldsymbol{\mu}(0)\cdot\boldsymbol{\mu}(t)\Big\rangle\,dt$ (19)
$R_{\mathrm{iso}}(\omega)\propto\omega^{2}\int_{-\infty}^{+\infty}e^{-i\omega t}\Big\langle\gamma(0)\cdot\gamma(t)\Big\rangle\,dt$ (20)
$R_{\mathrm{aniso}}(\omega)\propto\omega^{2}\int_{-\infty}^{+\infty}e^{-i\omega t}\Big\langle\mathrm{Tr}\!\big[\boldsymbol{\beta}(0)\cdot\boldsymbol{\beta}(t)\big]\Big\rangle\,dt$ (21)
where $\boldsymbol{\mu}$ denotes the dipole moment, $\boldsymbol{\alpha}$ the polarizability tensor, $\omega$ the vibrational frequency, and the angular brackets represent an ensemble average over time origins. The polarizability tensor $\boldsymbol{\alpha}$ can be decomposed into an isotropic trace component and a fully symmetric anisotropic component, expressed as $\boldsymbol{\alpha}=\boldsymbol{\prescript{(2,0,1)}{}{\alpha}}+\boldsymbol{\prescript{(2,2,1)}{}{\alpha}}=\gamma\boldsymbol{I}+\boldsymbol{\beta}$, where $\gamma$ denotes the isotropic part and $\boldsymbol{\beta}$ represents the anisotropic part of the polarizability tensor.
External field
The material properties under an external electric field can be obtained through differential relations involving the polarization $\mathbf{P}$, Born effective charges $\mathbf{BEC}$, and the polarizability $\boldsymbol{\alpha}$. These quantities can be expressed as follows, where $\mathrm{U}$ denotes the electric enthalpy, $i$ is the atomic index, and $\mu$ and $\nu$ are Cartesian indices:
$\mathbf{P}_{\mu}=-\frac{\partial\mathrm{U}}{\partial\mathbf{E}_{\mu}}$ (22)
$\boldsymbol{\alpha}_{\mu\nu}=\frac{\partial\mathbf{P}_{\mu}}{\partial\mathbf{E}_{\nu}}$ (23)
$\mathbf{BEC}_{i\mu\nu}=\frac{1}{e}\left.\frac{\partial\mathbf{P}_{\mu}}{\partial\mathbf{r}_{i\nu}}\right|_{\mathbf{E}_{\mu}=0}$ (24)
It should be noted that in periodic systems the polarization is multivalued. To handle this, we follow the same strategy as ref. [60], adopting the minimum image convention to eliminate ambiguities due to polarization quanta in the loss function and evaluation metrics, thereby avoiding additional data preprocessing.
@@
All datasets and input files of this study will be made available on GitHub https://github.com/xvzemin/tace-fit upon publication.
@@
The implementation of TACE will be made available on GitHub https://github.com/xvzemin/tace upon publication.

### v2

@@ 2.11 Advantages and Limitations of ACE — REICO-AgPdCHO
MatPES-PBE
M3GNet [78] CHGNet [^1] CHGNet includes magmoms information. So there will be better performance in terms of energy. [77] TensorNet [14] TACE TACE [^3] The ellipsis indicates that we can flexibly adjust the number of parameters by changing the size of the hidden channels and by choosing whether or not to include element dependence, ranging from 0.98M to 29.8M. TACE
- - - Element-Agnostic [^2] We follow MACE in the naming of model modules. “Element-Agnostic” indicates that the parameters of this module (product basis here) do not depend on the elements. The results of MACE-MATPES-PBE-0 are not included here, as the model was pretrained on OMat24 and subsequently finetuned, with the MatPES dataset further subjected to an additional filtering step. [10] Element-Aware
E ↓ 45 31 36 35 31
F ↓ 181 136 148 105 101
S ↓ 0.888 0.642 0.700 0.655 0.607
d ↓ 0.42 0.43 0.37 0.39 0.38
$\mathrm{E_{f}}$ 110 82 81 80 72
$\mathrm{K_{VRH}}$ 26 24 18 17 16
$\mathrm{G_{VRH}}$ 25 21 15 15 13
$\mathrm{C_{v}}$ 27 23 13 12 11
$\mathrm{\frac{F}{F_{DFT}}}$ 0.97 0.91 0.93 0.93 0.94
Parameters 0.66M 2.70M 0.84M 0.98M 29.8M
@@ 3 Discussion
In our view, the most effective form of equivariance arises from tensor product operations (irreducible Cartesian tensor products/contractions, SO(3)/SO(2) spherical tensor products, and a series of other derived methods developed to accelerate computations). For the work presented in this paper, we believe that the effectiveness and accuracy of Cartesian models have already been well demonstrated. However, we also see current limitations in the implementation of irreducible Cartesian tensor products/contractions, and hereby discuss the future road-map.
Regarding the irreducible Cartesian tensor products/contractions, present implementation typically forms tensor products over all valid combinations ($\nu_{1},\nu_{2},\nu_{o}$, not path here) and then extracts the highest-weight irreducible component via ICTD. In contrast, a single path spherical tensor product can generate multiple irreducible components; importantly, Cartesian tensors can also yield multiple irreducible components from a single tensor product/contraction. As illustrated in SI Fig. S3, the $3^{\nu}$ elements in a Cartesian tensor are not entirely redundant; they allow other irreducible components with different weights to be generated from a single tensor product/contraction. The specific implementation approach has already been outlined in Fig. 1a. Consider a rank-$\nu$ Cartesian tensor with weight $\mathrm{\nu-1}$. Such a tensor has $3^{\nu}$ components, but only $\mathrm{2\nu-1}$ of them are useful. In principle, one can project with ICTD, map to spherical irreps (retaining the $\mathrm{2\nu-1}$ nonzeros), then reassemble a rank-$\mathrm{(\nu-1)}$ Cartesian tensor of weight $\mathrm{(\nu-1)}$. Although we have not yet implement this pathway, we will improve this part based on e3nn in the future and anticipate this approach to embed richer radial information, reduce path count, thus, improve both accuracy and speed.
Beyond the ideal route above, we have employed ICTD in two ways: the first is “max”, which uses only the highest weight for a given rank; the second is “all”, which utilizes all available weights. For the “max” approach, we do not elaborate further here. In the “all” case, we do not directly embed radial weights along the path; however, each path is still modulated by a learnable weight initialized to 1.0. Instead, after tensor product, the messages obtained on each edge are processed via ICTD, at which point radial information is embedded through $\sum_{\ell,q}R_{\ell,q}\cdot\prescript{{(\nu;\ell;q)}}{}{\mathbf{T}}$. In other words, the number of radial weights depends on the rank of the tensor rather than the possibile path. The “all” strategy does not necessarily yield better performance, as it is not our optimal implementation. Nonetheless, pairing a weight $\ell$ with a learnable radial weight is a natural and straightforward approach. Finally, we emphasize that our tensor product operations do not require the input tensors to be irreducible, reducible cases can be handled in a unified manner.
Further speedups are available via compilation and kernel libraries. In current benchmarks (see SI 2.1), TACE is about half as fast as Cueq-accelerated MACE(v0.3.14). The current Cartesian framework does have ample potentials: removing the product basis and using layers with $\mathrm{T}>2$ (plus optional gating, i.e., a NequIP-like Cartesian architecture) yields significant speedups; even without acceleration, TACE ($\mathrm{T}>2$ , correction=1) is comparable to compiled NequIP(v0.15.0), while retaining an atomic basis and the functionality discussed above.
SO(2)-style simplification in Cartesian space. It is known that SO(2) spherical tensor products significantly simplify computational cost, making $\mathrm{L_{max}}$ or $\mathrm{\ell_{max}}$ = 6 feasible. In Cartesian space, it may be advantageous to use symmetric edge attributes. For example, if we rotate all edges to the direction $[1,0,0]$ (i.e., along the $x$-axis), then fully symmetric edge attributes with a trace will have only a single nonzero element. This implies that tensor products could effectively reduce to simple indexing and scaling operations. However, in our experiments, we found that the speed did not significantly improve, since contraction steps with rotation matrices are required for both the node features and the output features. Moreover, because the edge attributes used here are symmetric and not strictly equivalent to the SO(2) spherical tensor products, further improvement may be needed in the future.
$(\mathbf{E_{ij}})_{i_{1}i_{2}\dots i_{n}}=\frac{(2\nu_{2}-1)!!}{\nu_{2}!}\cdot\prod_{k=1}^{\nu_{2}}\delta_{i_{k},0},$ (20)
The gradient of the Cartesian tensor product is simple. Since the Cartesian tensor product does not involve Clebsch-Gordan coefficients, the tensor product can be reduced to simple matrix multiplication. Therefore, its gradient is very simple. Denoting the upstream gradient as G, the gradients with respect to inpute A and B during backpropagation are given as in Eqn. 21. By extension, arbitrary-order derivatives can be obtained. However, in our experiments, explicitly writing out the first analytical gradient did not lead to a speedup. This may be because PyTorch’s native optimizations are already efficient, or because our implementation still operates entirely within PyTorch without leveraging other backends. Nevertheless, it is possible that future developments could theoretically take better advantage of this property.
$\nabla_{A}=GB^{\top}~~~~~~\nabla_{B}=A^{\top}G$ (21)
Hybrid designs are also viable, for example, applying Cartesian tensor products with spherical representations. In the future, we will also introduce Cartesian-3j and Cartesian-nj to explore some new possibilities.
@@ 4 Conclusion
In this work, we present a unified design approach for Cartesian models and, based on this irreducible Cartesian tensor product/contracion method, implement TACE, an E(3)-equivariant (or SE(3)) message-passing framework built upon higher-order Cartesian tensors that operates directly in Cartesian space. TACE accommodates tensors of arbitrary rank and both reducible or irreducible Cartesian tensors. The framework can consistently extract the corresponding irreducible components of different weight, allowing us to freely extract and manipulate them. Furthermore, the model is capable of directly predicting tensorial properties, and through the use of a universal embedding, we can flexibly incorporate and control various factors (both invariants and equivariants) influencing the predicted physical quantities. Although we only demonstrate external fields, magnetic moments, charges, and fidelities here, the framework can be readily extended. Taken together, these results suggest that a Cartesian framework such as TACE can serve as the backbone of next-generation MLIPs, applicable across a broad range of problems in chemistry and physics and it is reasonable to expect that MLIPs of this style will become the primary computational engine for routine atomistic simulation and property prediction.
@@ 5 Methods
All experiments were conducted using the TACE software available at https://github.com/xvzemin/tace and Python version 3.12.11. The TACE-MatPES model was trained using PyTorch 2.8.0+cu128 and PyTorch Lightning 2.5.5, while all other results were obtained based on PyTorch 2.6.0+cu124 and PyTorch Lightning 2.5.2. For other models, we use orb-models 0.5.2, mace-torch 0.3.12(uMLIP), mace-torch 0.3.14(urea and speed test), cuequivariance 0.3.0, sevenn 0.11.1.dev2, allegro-pol 0.0.3 and nequip 0.15.0. In addition, most of the plotting relies on Matplotlib [79].
Bechmarks
@@
Water: For the liquid water dataset [46], a cutoff of 6 Å is applied. Following the protocol used in CAMP [17], the dataset is split into 1,432 configurations for training, 79 for validation, and 81 for testing. The dipole moment ($\mu$) and polarizability ($\alpha$) for various water systems were taken from ref. [44]. Values of $\mu$ and $\alpha$ for molecules such as $\mathrm{H_{2}O}$, $(\mathrm{H_{2}O})_{2}$, and $\mathrm{H_{5}O_{2}}^{+}$ were calculated using the coupled cluster singles and doubles (CCSD) method with the d-aug-cc-pVTZ basis set. For liquid water, the dataset was generated using the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional together with ultra-soft pseudopotentials (USPPs). A cutoff of 6 Å was employed for training the dipole and polarizability models, except for $(\mathrm{H_{2}O})_{2}$, for which a cutoff of 10 Å was used. The training and validation splits were either 500:500 or 700:300. The dataset employed for IR and Raman simulations comprises 1,888 structures calculated using the strongly constrained and appropriately normed (SCAN) functional. Our model attains a RMSE of 2.43meV/atom for energy, 30.97meV/Å for forces, and 1.00 meV/atom for virials. The train-test split follows the TNEP scheme [45], with 1,510 structures used for training and 378 for testing, and an additional 10% of the training set reserved for validation. The MD system simulated consists of 512 water molecules, corresponding to a density of 1 $\mathrm{g/cm^{3}}$. A time step of 1 fs was employed, and temperature control is maintained using a Nosé-Hoover thermostat.
@@
Charges: The datasets for different charge states and charge-transfer systems [38] were generated using the PBE functional with a light computational setting. Spin-polarized calculations were performed for the Au2-MgO, NaCl, and Ag3 systems. The Ag, NaCl, carbon-chain, and Au2-MgO datasets contain 11,013, 5,000, 10,019, and 5,000 structures, respectively. For each dataset, 10% of the configurations were used as the validation set, and all reported errors are evaluated on the validation set. For the charge embedding benchmarks of the Ag, NaCl, carbon-chain, and Au2-MgO systems, the cutoff radii are 5.29, 5.29, 4.23, and 5.5 Å, respectively. When using charges together with energy and forces as training labels, the corresponding cutoff values are 6.0, 6.0, 4.23, and 6.0 Å, respectively.
@@
PdAgCHO: PdAgCHO dataset was generated using the RECIO method [68] and contains a total of 116,516 structures. For fine-tuning, 100,000 structures were randomly selected, while the remaining configurations were used for validation. For from-scratch training, we use all structures, and the learning rate scheduler adopts warmup followed by cosine annealing. All structures were computed at the PBE level. The fine-tuning models and TACE models was carried out by us with a uniform cutoff of 6 Å. It should be noted that the NequIP [8] model was not trained by us, and that the dataset partitioning scheme employed in NequIP differs from ours.
MatPES: The MatPES dataset [76] was generated for uMLIPs, with calculations performed under both PBE and r2SCAN functionals. In this work, we only use the PBE subset and adopt the same data splitting scheme as in the original publication.
@@
Although multiple random seeds were not used in this study for efficiency, most tests were performed with seed = 42. We ensure that all results are fully reproducible. All datasets will be made available on GitHub https://github.com/xvzemin/tace-fit.
@@
The implementation of TACE will be made available on GitHub https://github.com/xvzemin/tace and uMLIPs training on OMat24 and several other datasets will also be released soon.